# FROST-CLUSTERS – I. Hierarchical star cluster assembly boosts intermediate-mass black hole formation

### v1

@@ 4 IMBHs from star cluster assembly
Cluster IMBH $M_{\mathrm{\star}}$ $t_{\mathrm{acc}}$ $M_{\mathrm{\bullet}}(t_{\mathrm{acc}})$
$\mathrm{[M_{\odot}]}$ $\mathrm{[Myr]}$ $\mathrm{[M_{\odot}]}$
H1-A $2.0\times 10^{5}$
H1-B BH1-B $4.4\times 10^{4}$ 6.5 $1194$
H1-C BH1-C $1.1\times 10^{5}$ 3.1 $536$
H1-D BH1-D $5.9\times 10^{4}$ 2.3 $348$
H2-A $2.3\times 10^{5}$
H2-B BH2-B $3.2\times 10^{4}$ 4.8 $1256$
H2-C BH2-C $2.6\times 10^{4}$ $350$
H3-A $2.3\times 10^{5}$
H3-B BH3-B $1.1\times 10^{5}$ 2.2 $1748$
H3-C BH3-C $1.5\times 10^{4}$ 8.7 $347$
H3-D BH3-D $2.3\times 10^{4}$ 3.7 $338$
H3-E BH3-E $1.4\times 10^{4}$ 6.9 $289$
H3-F BH3-F $2.6\times 10^{4}$ 7.6 $105$
Table 4: The initial stellar masses $M_{\mathrm{\star}}$ of the star clusters which produce an IMBH during the cluster assembly simulations. The most massive, central initial cluster is indicated with a label A. The accretion time $t_{\mathrm{acc}}$ indicates when the star cluster merges with the central cluster, and $M_{\mathrm{\bullet}}(t_{\mathrm{acc}})$ describes the mass of the most massive IMBH at this time. If $t_{\mathrm{acc}}$ is small enough, IMBHs have had no time to form yet. In this case the mass of the IMBH descending from the most massive star at the time of the accretion is given.
@@ 4.2 Merger trees of star clusters, their VMSs and IMBHs
The initial masses of the star clusters hosting IMBH progenitors are presented in Table 4. The table also presents the times $t_{\mathrm{acc}}$ when the clusters merge with the main central star cluster, and the IMBH mass at that time. If the IMBH has not formed yet at $t=t_{\mathrm{acc}}$, we provide the mass of the IMBH descending from the most massive VMS of the cluster at the time of its collapse.
@@ 4.3 Three categories of early IMBH growth histories
Table 5: The three categories of IMBHs formed in the hierarchical star cluster assembly simulations and their basic formation pathways. The IMBHs formed in the recent direct-summation N-body simulation studies of massive, isolated star clusters of Rizzuto et al. (2021); Rizzuto et al. (2022) and Arca Sedda et al. (2023c) would correspond the categories IMBH-LM and IMBH-JAM of this classification.
@@ 5 IMBHs from isolated star clusters
Cluster ($N_{\mathrm{seed}}$) $N$ $M_{\mathrm{\star}}$ $\max(m_{\star})$ $\max(m_{\bullet})$
$\mathrm{[M_{\odot}]}$ $\mathrm{[M_{\odot}]}$ $\mathrm{[M_{\odot}]}$
I1 ($10$) $1.7\times 10^{4}$ $9.4\times 10^{3}$ 197197{197}197 7272{72}72
I2 ($10$) $2.6\times 10^{4}$ $1.5\times 10^{4}$ 290290{290}290 8787{87}87
I3 ($10$) $3.9\times 10^{4}$ $2.3\times 10^{4}$ 419419{419}419 378378{378}378
I4 ($10$) $6.3\times 10^{4}$ $3.7\times 10^{4}$ 932932{932}932 964964{964}964
I5 ($10$) $1.0\times 10^{5}$ $5.9\times 10^{4}$ 982982{982}982 10651065{1065}1065
I6 ($10$) $1.6\times 10^{5}$ $9.4\times 10^{4}$ 14291429{1429}1429 15101510{1510}1510
I7 ($10$) $2.5\times 10^{5}$ $1.5\times 10^{5}$ 14941494{1494}1494 16721672{1672}1672
I8 ($20$) $4.0\times 10^{5}$ $2.3\times 10^{5}$ 362362{362}362 9191{91}91
I9 ($5$) $6.2\times 10^{5}$ $3.7\times 10^{5}$ 199199{199}199 8181{81}81
I10 ($5$) $1.0\times 10^{6}$ $5.9\times 10^{5}$ 150150{150}150 6363{63}63
@@
Table 7: The ten isolated star cluster merger sets I1–I10, in total $100$ isolated simulations with their initial particle numbers $N$, initial masses $M_{\star}$ and the masses for the most massive star ($\max{m_{\star}}$) and (IM)BH $(\max{m_{\bullet}}$) formed through collisions in the simulation sets.
@@
Figure 10: The maximum stellar (left panel) and the BH (right panel) masses $\max{(m_{\star})}$ and $\max{(m_{\bullet})}$ reached in the isolated star clusters during the first $t=5$ Myr of the simulations. The solid black line indicates the expectation without mergers from the Kroupa (2001) initial mass function a maximum mass of $m_{\star}={150}\>\mathrm{M_{\odot}}$ including the suppression of high-mass stars in low-mass clusters (Yan et al., 2023). The maximum final individual stellar mass grows linearly as $\max{(m_{\star})}\approx 0.02\times M_{\star}$ up to star cluster masses of $M_{\star}\sim{10^{5}}\>\mathrm{M_{\odot}}$, after which the VMS growth in more massive clusters is strongly suppressed. The maximum BH mass follows the trends of the maximum stellar mass, processed by stellar evolution. Below cluster masses $M_{\mathrm{\star}}\lesssim{2\times 10^{4}}\>\mathrm{M_{\odot}}$, no BHs with masses above the (P)PISN mass gap (in gray shading) form. The maximum reached IMBH masses are $m_{\bullet}\sim{1630}\>\mathrm{M_{\odot}}$. At cluster masses higher than $M_{\star}\gtrsim{2\times 10^{5}}\>\mathrm{M_{\odot}}$ there are no IMBHs above the (P)PISN mass, but BHs in the mass gap below the IMBH limit $m_{\bullet}={100}\>\mathrm{M_{\odot}}$ still form. Note that the maximum IMBH mass can exceed the maximum VMS mass as the IMBHs still grow via disrupting stars in the remnant phase.
@@ 5.6 Suppressing collisions between massive stars in massive star clusters
The velocity dispersion argument is consistent with the earlier findings in this section, especially the facts that mergers between massive stars from bound orbits are suppressed at sufficiently high cluster masses $M_{\star}$, and if collisions involving massive stars take place, they occur on average at later times. Finally, the life-times of the most massive stars ($t=2.82$ Myr for $m_{\star}={150}\>\mathrm{M_{\odot}}$) ensure that the collisions are not just postponed, but completely inhibited as the massive stars reach the ends the their lives before they manage to collide with each other.
@@ 6.4 Gravitational waves
The most numerous gravitational-wave merger category during the first $t=50$ Myr of the cluster assembly simulations in this study are IMBH-BH mergers. Most IMBH-BH mergers have their total masses $m_{\mathrm{b}}=m_{\mathrm{p}}+m_{\mathrm{s}}$ and mass ratios $q=m_{\mathrm{s}}/m_{\mathrm{p}}$ ranging from ${1300}\>\mathrm{M_{\odot}}\lesssim m_{\mathrm{b}}\lesssim{2220}\>\mathrm{M_{\odot}}$ and $1.5\times 10^{-2}\lesssim q\lesssim 4.5\times 10^{-2}$. One IMBH-BH merger has a lower total mass of $m_{\mathrm{b}}\sim{600}\>\mathrm{M_{\odot}}$ and $q\sim 0.1$. Finally, the only true IMBH-IMBH merger of this study has a total mass of $m_{\mathrm{b}}\sim{1500}\>\mathrm{M_{\odot}}$ and $q\sim 0.3$. Based on the literature work (e.g. Rizzuto et al. 2021; Rizzuto et al. 2022), we expect that the inclusion of primordial binary stars in the initial cluster setups in our forthcoming study FROST-CLUSTERS II will populate the BH-BH LIGO/Virgo/KAGRA parameter space in Fig. 16. Beside this regime, our hierarchical star cluster assembly simulations also indicate a distinct second GW population in the IMBH mass range around $m_{\mathrm{b}}\sim{10^{3}}\>\mathrm{M_{\odot}}$ with typical mass ratios of $1.5\times 10^{-2}\lesssim q\lesssim 4.5\times 10^{-2}$.
@@ 7 Summary and conclusions
Our main findings are the following. First, IMBHs form in each three hierarchical cluster assembly simulations during the first $t=50$ Myr with maximum IMBH masses $m_{\bullet}\sim{2200}\>\mathrm{M_{\odot}}$. The primary formation channel for IMBH formation is the collapse of a VMS which formed through a sequence of collisions with massive stars, the primary progenitor having an initial mass exceeding $m_{\star}\gtrsim{80}\>\mathrm{M_{\odot}}$. Tidal disruption of stars, BH-BH, IMBH-BH and IMBH-IMBH (in one case) mergers also contribute to the final mass of the IMBHs, but their contribution is sub-dominant during the first $50$ Myr. Typically only a single IMBH forms in a star cluster, but multiple IMBHs will end up in the most massive star cluster through the cluster assembly process. Multiple IMBHs per final star cluster is essentially an unique prediction of the model as in isolated setups typically no more than a single object can considerably grow by stellar mergers.
@@
The suppression of VMS growth and rapid IMBH formation in massive, isolated star clusters caused by a combination of two factors. In star clusters with masses in the range of ${2\times 10^{4}}\>\mathrm{M_{\odot}}\lesssim M_{\star}\lesssim{2\times 10^{5}}\>\mathrm{M_{\odot}}$ binaries with massive stars can form through the 3BBF channel and subsequent dynamical three-body exchanges. As the velocity dispersion of these clusters is relatively low ($\lesssim 40$ km/s), these binaries can harden according to Heggie-Hills law, and combined with their high eccentricities and low pericenter distances also eventually merge with each other. In the case of more massive clusters the velocity dispersion is high and the 3BBF channel is suppressed. If binaries with massive stars form, they are too wide to harden with further interactions with the cluster stars. Thus, massive stellar mergers and the VMS growth is inhibited. Second, the relatively short life-times of massive stars ($t=2.82$ Myr for $m_{\star}={150}\>\mathrm{M_{\odot}}$) ensure that the onset of massive star mergers is truly prevented and not just postponed as the massive stars reach the ends of their lives.

### v2

@@ 4 IMBHs from star cluster assembly
Cluster IMBH $M_{\mathrm{\star}}$ $\rho_{\mathrm{c,init}}$ $t_{\mathrm{seg}}$ $t_{\mathrm{cc}}$ $t_{\mathrm{acc}}$ $t_{\mathrm{cc}}<t_{\mathrm{acc}}?$ $M_{\mathrm{\bullet}}(t_{\mathrm{acc}})$
$\mathrm{[M_{\odot}]}$ $[{10^{6}}\>\mathrm{M_{\odot}\>pc^{-3}}]$ $\mathrm{[Myr]}$ $\mathrm{[Myr]}$ $\mathrm{[Myr]}$ $\mathrm{[M_{\odot}]}$
H1-A $2.0\times 10^{5}$ 1.77 0.32 7.60
H1-B BH1-B $4.4\times 10^{4}$ 0.88 0.10 1.95 6.5 $\checkmark$ $1194$
H1-C BH1-C $1.1\times 10^{5}$ 1.34 0.20 4.43 3.1 - $536$
H1-D BH1-D $5.9\times 10^{4}$ 1.01 0.12 2.53 2.3 - $348$
H2-A $2.3\times 10^{5}$ 1.89 0.35 8.63
H2-B BH2-B $3.2\times 10^{4}$ 0.76 0.08 1.47 4.8 $\checkmark$ $1256$
H2-C BH2-C $2.6\times 10^{4}$ 0.69 0.08 1.23 - $350$
H3-A $2.3\times 10^{5}$ 1.89 0.35 8.63
H3-B BH3-B $1.1\times 10^{5}$ 1.34 0.20 4.43 2.2 - $1748$
H3-C BH3-C $1.5\times 10^{4}$ 0.54 0.06 0.76 8.7 $\checkmark$ $347$
H3-D BH3-D $2.3\times 10^{4}$ 0.65 0.07 1.10 3.7 $\checkmark$ $338$
H3-E BH3-E $1.4\times 10^{4}$ 0.52 0.06 0.71 6.9 $\checkmark$ $289$
H3-F BH3-F $2.6\times 10^{4}$ 0.69 0.08 1.23 7.6 $\checkmark$ $105$
Table 4: The initial stellar masses $M_{\mathrm{\star}}$ of the star clusters which produce an IMBH during the cluster assembly simulations. The most massive, central initial cluster is indicated with a label A. The estimates for the mass segregation and core collapse time-scales ($t_{\mathrm{seg}}$, $t_{\mathrm{cc}}$) are obtained using Eq. (3) and (4).
The accretion time $t_{\mathrm{acc}}$ indicates when the star cluster merges with the central cluster, and $M_{\mathrm{\bullet}}(t_{\mathrm{acc}})$ describes the mass of the most massive IMBH at this time. If $t_{\mathrm{acc}}$ is small enough, IMBHs have had no time to form yet. In this case the mass of the IMBH descending from the most massive star at the time of the accretion is given.
@@ 4.2 Merger trees of star clusters, their VMSs and IMBHs
The initial masses $M_{\star}$ and central densities $\rho_{\mathrm{c,init}}$ of the star clusters hosting IMBH progenitors are presented in Table 4. The table also presents the times $t_{\mathrm{acc}}$ when the clusters merge with the main central star cluster, and the IMBH mass at that time. If the IMBH has not formed yet at $t=t_{\mathrm{acc}}$, we provide the mass of the IMBH descending from the most massive VMS of the cluster at the time of its collapse. The table also shows the estimates for the mass segregation and core collapse time-scales ($t_{\mathrm{seg}}$, $t_{\mathrm{cc}}$) for the clusters based on Eq. (3) and Eq. (4) with $t_{\mathrm{cc}}\sim 0.2\;t_{\mathrm{rlx}}$ (Portegies Zwart et al., 2004). Based on the estimate, mass segregation occurs rapidly in all models with $t_{\mathrm{seg}}<t_{\mathrm{acc}}$. However, not all clusters that form IMBHs are core collapsed when they merge with the central cluster. Thus, core collapse seems not to be the absolute bottleneck for collisional VMS and IMBH formation in the simulations of this study. Dense, mass segregated clusters can support repeated stellar collisions already at times $t_{\mathrm{seg}}<t<t_{\mathrm{cc}}$.
@@ 4.3 Three categories of early IMBH growth histories
Table 5: The three categories of IMBHs formed in the hierarchical star cluster assembly simulations and their basic formation pathways. The IMBHs formed in the recent direct-summation N-body simulation studies of massive, isolated star clusters of Rizzuto et al. (2021); Rizzuto et al. (2022) and Arca Sedda et al. (2023) would correspond the categories IMBH-LM and IMBH-JAM of this classification.
@@ 5 IMBHs from isolated star clusters
Cluster $N_{\mathrm{seed}}$ $N$ $M_{\mathrm{\star}}$ $\rho_{\mathrm{c,init}}$ $t_{\mathrm{seg}}$ $t_{\mathrm{cc}}$ $\max(m_{\star})$ $\max(m_{\bullet})$
$\mathrm{[M_{\odot}]}$ ${10^{6}}\>\mathrm{M_{\odot}\>pc^{-3}}$ Myr Myr $\mathrm{[M_{\odot}]}$ $\mathrm{[M_{\odot}]}$
I1 $10$ $1.7\times 10^{4}$ $9.4\times 10^{3}$ 0.43 0.05 0.41 197197{197}197 7272{72}72
I2 $10$ $2.6\times 10^{4}$ $1.5\times 10^{4}$ 0.54 0.05 0.62 290290{290}290 8787{87}87
I3 $10$ $3.9\times 10^{4}$ $2.3\times 10^{4}$ 0.65 0.06 0.91 419419{419}419 378378{378}378
I4 $10$ $6.3\times 10^{4}$ $3.7\times 10^{4}$ 0.81 0.07 1.38 932932{932}932 964964{964}964
I5 $10$ $1.0\times 10^{5}$ $5.9\times 10^{4}$ 1.00 0.10 2.09 982982{982}982 10651065{1065}1065
I6 $10$ $1.6\times 10^{5}$ $9.4\times 10^{4}$ 1.25 0.14 3.17 14291429{1429}1429 15101510{1510}1510
I7 $10$ $2.5\times 10^{5}$ $1.5\times 10^{5}$ 1.55 0.21 4.83 14941494{1494}1494 16721672{1672}1672
I8 $20$ $4.0\times 10^{5}$ $2.3\times 10^{5}$ 1.89 0.29 7.11 362362{362}362 9191{91}91
I9 $5$ $6.2\times 10^{5}$ $3.7\times 10^{5}$ 2.34 0.43 10.95 199199{199}199 8181{81}81
I10 $5$ $1.0\times 10^{6}$ $5.9\times 10^{5}$ 2.91 0.63 16.76 150150{150}150 6363{63}63
@@
Table 7: The ten isolated star cluster simulation sets I1–I10, in total $100$ isolated simulations with their initial particle numbers $N$, initial masses $M_{\star}$ and the masses for the most massive star ($\max{m_{\star}}$) and (IM)BH $(\max{m_{\bullet}}$) formed through collisions in the simulation sets. We also show the initial central densities of the cluster models $\rho_{\mathrm{c,init}}$ as well as estimates for the mass segregation and core collapse time-scales ($t_{\mathrm{seg}}$, $t_{\mathrm{cc}}$) of the models as in Table 4.
@@
Figure 10: The maximum stellar (left panel) and the BH (right panel) masses $\max{(m_{\star})}$ and $\max{(m_{\bullet})}$ reached in the isolated star clusters during the first $t=5$ Myr of the simulations. The solid black line indicates the expectation without mergers from the Kroupa (2001) initial mass function a maximum mass of $m_{\star}={150}\>\mathrm{M_{\odot}}$ including the suppression of high-mass stars in low-mass clusters (Yan et al., 2023). The maximum final individual stellar mass grows linearly as $\max{(m_{\star})}\approx 0.02\times M_{\star}$ up to star cluster masses of $M_{\star}\sim{10^{5}}\>\mathrm{M_{\odot}}$, after which the VMS growth in more massive clusters is strongly suppressed. The maximum BH mass in the isolated runs (colored crosses) follows the trends of the maximum stellar mass, processed by stellar evolution. Below cluster masses $M_{\mathrm{\star}}\lesssim{2\times 10^{4}}\>\mathrm{M_{\odot}}$, no BHs with masses above the (P)PISN mass gap (in gray shading) form. The maximum reached IMBH masses are $m_{\bullet}\sim{1630}\>\mathrm{M_{\odot}}$. At cluster masses higher than $M_{\star}\gtrsim{2\times 10^{5}}\>\mathrm{M_{\odot}}$ there are no IMBHs above the (P)PISN mass, but BHs in the mass gap below the IMBH limit $m_{\bullet}={100}\>\mathrm{M_{\odot}}$ still form. Note that the maximum IMBH mass can exceed the maximum VMS mass as the IMBHs still grow via disrupting stars in the remnant phase. For comparison, the IMBH (or their progenitor) masses from the hierarchical runs at the time of their accretion into the main cluster (as in Table 4) are marked with black, open circles. The IMBHs formed in the in-falling sub-clusters can be more massive than their counterparts in isolated star clusters of the same mass and initial density.
@@ 5.4 The time of onset of stellar mergers
The theoretical estimates for the two-body relaxation and mass segregation time-scales (as in Eq. (3) and Eq. (4)) monotonically increase with increasing cluster mass for the cluster mass-size relation adopted for this study. In Fig. 13 we see that in our simulations the first collisions occur earlier in more massive clusters. Compared to the core collapse time-scales (as determined from the Lagrangian radii) of the simulation models, the first collisions in the low-mass models occur considerably later than the core collapse time ($t_{\mathrm{cc}}\gtrsim 0.6$ Myr). We note that the core collapse time-scales for the massive star clusters appear to be by up to a factor of few shorter when determined from the Lagrangian radii instead of the analytical estimate of Table 7 based on the two-body relaxation time-scale. For higher-mass (I6-I10) models, the core collapse time is closer to the time of the first collision involving a high-mass star. In summary, no simple time-scale estimate explains the trends in Fig. 13 in a straightforward manner. It is likely that both the central density itself and the mass segregation and the core collapse time-scales all play a role in determining the times for the first collisions.
@@ 5.6 Suppressing collisions between massive stars in massive star clusters
The velocity dispersion argument is consistent with the earlier findings in this section, especially the facts that mergers between massive stars from bound orbits are suppressed at sufficiently high cluster masses $M_{\star}$, and if collisions involving massive stars take place, they occur on average at later times. Finally, the life-times of the most massive stars ($t=2.82$ Myr for $m_{\star}={150}\>\mathrm{M_{\odot}}$) ensure that the collisions are not just postponed, but completely inhibited as the massive stars reach the ends the their lives before they manage to collide with each other. Our results highlight the importance of the efficiency of the dynamical binary formation channel for the repeated stellar collisions in mass segregated but not yet core collapsed ($t_{\mathrm{seg}}<t<t_{\mathrm{cc}}$) massive star clusters.
@@ 6.4 Gravitational waves
The most numerous gravitational-wave merger category during the first $t=50$ Myr of the cluster assembly simulations in this study are IMBH-BH mergers. Most IMBH-BH mergers have their total masses $m_{\mathrm{b}}=m_{\mathrm{p}}+m_{\mathrm{s}}$ and mass ratios $q=m_{\mathrm{s}}/m_{\mathrm{p}}$ ranging from ${1300}\>\mathrm{M_{\odot}}\lesssim m_{\mathrm{b}}\lesssim{2220}\>\mathrm{M_{\odot}}$ and $1.5\times 10^{-2}\lesssim q\lesssim 4.5\times 10^{-2}$. One IMBH-BH merger has a lower total mass of $m_{\mathrm{b}}\sim{600}\>\mathrm{M_{\odot}}$ and $q\sim 0.1$. Finally, the only true IMBH-IMBH merger of this study has a total mass of $m_{\mathrm{b}}\sim{1500}\>\mathrm{M_{\odot}}$ and $q\sim 0.3$. Based on the literature work (e.g. Di Carlo et al. 2019; Rodriguez et al. 2019; Banerjee et al. 2020; Rastello et al. 2021; Rizzuto et al. 2022; Torniamenti et al. 2022; Arca Sedda et al. 2024b), we expect that the inclusion of primordial binary stars in the initial cluster setups in our forthcoming study FROST-CLUSTERS II will populate the BH-BH LIGO/Virgo/KAGRA parameter space in Fig. 16. Beside this regime, our hierarchical star cluster assembly simulations also indicate a distinct second GW population in the IMBH mass range around $m_{\mathrm{b}}\sim{10^{3}}\>\mathrm{M_{\odot}}$ with typical mass ratios of $1.5\times 10^{-2}\lesssim q\lesssim 4.5\times 10^{-2}$.
@@ 7 Summary and conclusions
Our main findings are the following. First, IMBHs form in each three hierarchical cluster assembly simulations during the first $t=50$ Myr with maximum IMBH masses $m_{\bullet}\sim{2200}\>\mathrm{M_{\odot}}$. The primary formation channel for IMBH formation is the collapse of a VMS which formed through a sequence of collisions with massive stars, the primary progenitor having an initial mass exceeding $m_{\star}\gtrsim{80}\>\mathrm{M_{\odot}}$. All star clusters forming IMBHs are mass segregated before they merge with the central cluster, but not all of them have reached the core collapse. Tidal disruption of stars, BH-BH, IMBH-BH and IMBH-IMBH (in one case) mergers also contribute to the final mass of the IMBHs, but their contribution is sub-dominant during the first $50$ Myr. Typically only a single IMBH forms in a star cluster, but multiple IMBHs will end up in the most massive star cluster through the cluster assembly process. Multiple IMBHs per final star cluster is essentially an unique prediction of the model as in isolated setups typically no more than a single object can considerably grow by stellar mergers.
@@
The suppression of VMS growth and rapid IMBH formation in massive, isolated star clusters caused by a combination of two factors. When massive stars end their lives, their dense, massive star clusters have mass segregated but not necessarily yet core collapsed. In star clusters with masses in the range of ${2\times 10^{4}}\>\mathrm{M_{\odot}}\lesssim M_{\star}\lesssim{2\times 10^{5}}\>\mathrm{M_{\odot}}$ binaries with massive stars can form through the 3BBF channel and subsequent dynamical three-body exchanges. As the velocity dispersion of these clusters is relatively low ($\lesssim 40$ km/s), these binaries can harden according to Heggie-Hills law, and combined with their high eccentricities and low pericenter distances also eventually merge with each other. In the case of more massive clusters the velocity dispersion is high and the 3BBF channel is suppressed. If binaries with massive stars form, they are too wide to harden with further interactions with the cluster stars. Thus, massive stellar mergers and the VMS growth is inhibited. Second, the relatively short life-times of massive stars ($t=2.82$ Myr for $m_{\star}={150}\>\mathrm{M_{\odot}}$) ensure that the onset of massive star mergers is truly prevented and not just postponed as the massive stars reach the ends of their lives. We also note that as the assumed star cluster mass-size relation (an observationally motivated power-law with $\beta=0.18$) is relatively shallow, so the relative increase in the central densities of the cluster models e.g. from $N=10^{5}$ to $N=10^{6}$ is only moderate. For more compact million-body models the core-collapse time-scale would be shorter, potentially allowing for more mergers of massive stars and thus IMBH formation.